Exercise 12-5 Volume Trade-Off Decisions [LO12-5]

Outdoor Luggage, Inc., makes high-end hard-sided luggage for sports equipment. Data concerning three of the company’s most popular models appear below.

Ski
Guard
Golf
Guard
Fishing
Guard
Selling price per unit $ 220 $ 330 $ 200
Variable cost per unit $ 140 $ 260 $ 115
Plastic injection molding machine processing
time required to produce one unit
4 minutes 12 minutes 9 minutes
Pounds of plastic pellets per unit 12 pounds 15 pounds 13 pounds

Required:

1. If we assume that the total time available on the plastic injection molding machine is the constraint in the production process, how much contribution margin per minute of the constrained resource is earned by each product?

2. Which product offers the most profitable use of the plastic injection molding machine?

3. If we assume that a severe shortage of plastic pellets has required the company to cut back its production so much that its new constraint has become the total available pounds of plastic pellets, how much contribution margin per pound of the constrained resource is earned by each product?

4. Which product offers the most profitable use of the plastic pellets?

5. Which product has the largest contribution margin per unit?
Answer #1
Ski
Guard
Golf
Guard
Fishing
Guard
Selling price per unit (i) 220 330 200
Variable cost per unit (ii) 140 260 115
Contribution margin per unit (iii) = (i) - (ii) 80 70 85
Plastic injection molding machine processing
time required to produce one unit (iv)
4 minutes
12 minutes
9 minutes
Contribution margin per minute of Plastic injection molding machine (iii)/(iv) 80/4 = $20 70/12 = $5.83 85/9 = $9.44
Pounds of contribution margin per pound per unit (v)
12 pounds
15 pounds 13 pounds
Contribution margin per pound of plastic pellets (iii)/(v) 80/12 = $6.67 70/15 = $4.67 85/13 = $6.54

1.

Ski
Guard
Golf
Guard
Fishing
Guard
Contribution margin per minute of Plastic injection molding machine $20 $5.83 $9.44

2.

Since Contribution margin per minute of Plastic injection molding machine of Ski Guard is highest, hence Ski
Guard offers the most profitable use of the plastic injection molding machine.

3.

Ski
Guard
Golf
Guard
Fishing
Guard
Contribution margin per pound of plastic pellets $6.67 $4.67 $6.54

4.

Since Contribution margin per pound of plastic pellets of Ski Guard is highest, hence Ski
Guard offers the most profitable use of the plastic pellets.

5.

Fishing Guard has the largest contribution margin per unit.

Kindly give a positive rating if you are satisfied with the answer. Feel free to ask if you have any doubt. Thanks